TEST DIAGNOSTICS.WORKING WITH INTERNAL LAN

CLOSE LOOP @ DIFFERENT FREQUENCIES CHANGING THE BUFFER DIMENSION.


aoserver buffer change to 64.
CLOSE LOOP 300HZ.
frames saved :10000
400Hz
500Hz.
Nota: analysis of the frame counter file, at 400 and 500Hz loose too many packets, 11 and 18 repectively. At 300Hz only has lost one packet.

Change the buffer dimension to 200.
500Hz.
nota: looks the same behaves than before, it is not improving.

Try to change the dimesion of the buffer in teh adsec. to 200. at 500Hz the results is the same.apply an update in tehy seconmdary.

nota: after applying open loop the wfs gui went to failure.
Stop testing procedure working on the fault.
TEST diagonstic speed with update ON.

IMPROVEMENTS IN THE ACQUISITION AT 400 AND 500 Hz.LOOSING FEW PACKETS (avg. 1 every 5000).


ACQUISITION OF SLOPES FOR THE SINUSOIDAL PROCEDURE.

INTERACTION MATRIX AT DIFFERENT BINNING.


bin:2.
TT:3.
speed:600Hz.


generate dist. 20090625_135905
bin:2
Nmodes:153.
ccd39: 600.
ampl:pp_m650_b2_ccd600_tt3.fits.
frames:5
cycle:2.


Rec: REC_20090625_140234
Nmodes:153.
ccd39:600Hz.
TT:3
Intmat generate: _ 140011 file slopes
Iter:4
Intmat: _140234

nota: rip the secondary.two times to apply the flatao succesful.
Error of the WFS control with the devices on and off continously.. Problem discovered...

Test of the Rec acquire.

clòose loop ..Ok.
close loop + atmos disturbance... the secondary start to skip frames and stop the secondary. Rip the secondary to recover the secondary.

Acquire a new rec after adjust the pupils. The number of subapertures 146.

New rec. REC_155336.fits
cut 7 modes.
Nmodes:153.
ccd39:600Hz.
TT:3
Intmat generate: _ 140011 file slopes
Iter:4
Intmat: _155336.

nota: close the loop but it is loosing to many modes in the border. The close loop didn't looks right.
Rip the secondary.

Try to change the mode number, let say 120.



REC_20090625_160812

-Acquire the Intmat.
CmdDistur_20090625_135905
-analyse the intmat.
cut 120.
reject 1 mode.

close loop
gain:0.6
nmodes:120.
ccd39:600Hz.
Bin:2
TT:3
Disturb: _1000Hz_

Nota: Close loop ok....

LOW FLUX test.

ccd39. 600hz
bin:2
Counts:80
FW1. 600-1000.
FW2.850-10.
-------------------------------------------------------------------------------------
saved psf files:
directory: /towerdata/adsec_data/Data_20090625/.
PSF.fits


Rec:REC_20090625_160812
nModes:120.
aveg:500.
ccd39. 600hz
bin:2
Counts:80
FW1. 600-1000.
FW2.850-10.
TT:3
Gain:0.6.
slopes rms:1.2e-1
PSF2.fits.

aveg:500.
same configuration.

NOTA: Changing the ccd47 speed to 0.4Hz.
PSF5.fits.

aveg. frame: 200
same configuration.
Optical loop data :data saved: Data_20090625_173957.

SUMMARY: We need to improve the pupil registration. put the ccd47 speed for low flux increase to 0.4Hz.



*************SINUSOIDAL MEASURES

Nota: mirror rip the last measurements 45Hz have to be acquire again.

-- JuanCarlosGuerra - 2009-06-26
Topic revision: r2 - 15 Jul 2009, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback